For the reaction



MnO2(gas) + 4 HCl(aq) ----------> MnCl2(aq) + Cl2(g) + 2 H2O(l)





1. what volume of Cl2(g) measured at STP is produced when 7.65 g of HCl(aq) reacts?

1
Expert's answer
2022-05-19T00:28:03-0400

MnO2 + 4 HCl = MnCl2 + Cl2 + 2 H2O

1. we find the substance quantity of HCl. To do this, we divide the mass of HCl by its molecular mass(36,5gr/mol) :

7.65 / 36,5 = 0.21 mol.

2. According to the above reaction, 1 mole of Cl2 is formed when 4 moles of HCl are reacted. Using this, we can find out how many moles of Cl2 are formed when 0.21 mol HCl reacts:

4mol HCl —> 1mol Cl2

0.21mol HCl –> x mol Cl2

X= 0.21 * 1 / 4 = 0.0525 mol

3. under normal conditions, 1 mole of gaseous substance has a volume of 22.4 liters. Using this information, we find the volume of 0.0525 mol of Cl2. To do this, we multiply the amount of substance (mol) by 22.4 litres:

V= 0.0525 * 22.4 = 1.176 litres.

ANSWER: 1.176 litres Cl2